# NPGA: Neural Parametric Gaussian Avatars

### v1

@@ 5.5. Ablations Study
In order to verify several important components of NPGA, we perform ablation experiments using three subjects. Quantitative and qualitative results of our ablations can be found in Table 2 and Fig. 5, respectively. First, we run a ”vanilla” version of NPGA that serves as a baseline. This version does not utilize per Gaussian features, Laplacian smoothness terms, and uses $e=1$ in Eq. 13 for the ADC. This model fails to produce sharp renderings for fine-scale details, and areas that are complicated due to frequent occlusions and reflections, like the bottom teeth and eyes.
@@
When adding per Gaussian features to the vanilla model, denoted as ”+p.G.F.”, the increased representational capacity results in sharper reconstructions. At the same time, we occasionally observe artifacts of ”free-floating” primitives, as highlighted in the second column of Fig. 5. These artifacts can be largely removed using our proposed Laplacian smoothness terms. Table 2 indicates that using this regularization significantly shrinks the generalization gap between training (NVS) and testing (self-reenactment) sequences. Compared to this model, ”Ours” also includes a CNN, which further boosts metrics and visual quality.
@@
Finally, we show the importance of using an adopted ADC strategy. Surprisingly, using the default ADC settings with a densification interval of 5000 steps and an opacity reset interval of $50.000$ steps almost completely diminishes the improvements of our other contributions. While we do not claim that using $e=2$ in Eq. 13 is necessary for great performance, we simply note that finding a setting that lets enough Gaussians appear is important, especially for fine-scaled wrinkles and teeth.
@@ 7. Conclusion
In this work, we have proposed Neural Parametric Gaussian Avatars (NPGA), a method for creating accurately controllable and high-fidelity virtual head avatars. A key component of our work is the usage of MonoNPHM’s rich expression space and motion prior. To this end, we propose a cycle-consistency strategy to distill a forward deformation field from MonoNPHM, such that it becomes compatible with 3D Gaussian Splatting. Furthermore, we have introduced per Gaussian features, a simple technique, to help our deformation modules explain the dynamic behavior of our avatar’s canonical Gaussian primitives. We proposed an effective regularization strategy and adjusted the adaptive density control strategy, both of which are important for ideal avatar quality. In our experiments, we significantly outperform the previous state-of-the-art avatars on self-reenactment. Finally, we showed the applicability of our avatars beyond a controlled multi-view set-up by animating them from monocular RGB video trackings.

### v2

@@ 5.5. Ablations Study
In order to verify several important components of NPGA, we perform ablation experiments using three subjects. Quantitative and qualitative results of our ablations can be found in Table 2 and Fig. 5, respectively. First, we run a ”vanilla” version of NPGA that serves as a baseline. This version does not utilize per Gaussian features, the CNN, Laplacian smoothness terms, and uses $e=1$ in Eq. 13 for the ADC. This model fails to produce sharp renderings for fine-scale details, and areas that are complicated due to frequent occlusions and reflections, like the bottom teeth and eyes.
@@
When adding per-Gaussian features to the vanilla model, denoted as ”+p.G.F.”, the increased representational capacity results in sharper reconstructions. At the same time, we occasionally observe artifacts of ”free-floating” primitives, as highlighted in the second column of Fig. 5. These artifacts can be largely removed using our proposed Laplacian smoothness terms. Table 2 indicates that using this regularization significantly shrinks the generalization gap between training (NVS) and testing (self-reenactment) sequences. Compared to this model, ”Ours” also includes a CNN, which further boosts metrics and visual quality. Additionally, we ablate the importance of letting per-Gaussian features influence the predicting of position offsets, since recent work (Xu et al., 2024; Dhamo et al., 2023) does not allow such interactions. In Table 2 ”Ours-$\delta_{\mu}$” denotes a version of NPGA that disables such an influence, indicating the benefits of our formulation.
@@
Finally, we show the importance of using an adjusted ADC strategy. Surprisingly, using the default ADC settings with a densification interval of 5000 steps and an opacity reset interval of $50.000$ steps almost completely diminishes the improvements of our other contributions. While we do not claim that using $e=2$ in Eq. 13 is necessary for great performance, we simply note that finding a setting that lets enough Gaussians appear is important, especially for fine-scaled wrinkles and teeth. The progress of the number of Gaussian during the optimization is illustrated in Fig. 8.
Figure 7. Tracking Failure: While our MonoNPHM-based tracking occasionally fails for extreme facial expressions (see normals (i) and overlay (ii)), NPGA still produces good animations for held-out expressions (see (iv)) due to the smooth nature of expression latent space.
Figure 8. Number of Gaussians: We illustrate the growing number of Gaussians for each of our avatars, where PID denotes the personal identifiers in the Nersemble dataset. We limit the maximum number to 250k.
@@ 6. Limitations and Future Work
7. Ethical Considerations
The creation of photo-realistic avatars has potential for a wide range of malevolent activities, e.g. privacy violation, identity theft and spreading of deceptive content through deepfakes. We condemn any malicious or unauthorized uses of such technology, and emphasize the need for reliable detection methods that maintain the authenticity of media content and advert negative societal impact.
@@ 8. Conclusion
In this work, we have proposed Neural Parametric Gaussian Avatars (NPGA), a method for creating accurately controllable and high-fidelity virtual head avatars. The main focus of our work is the usage of MonoNPHM’s rich expression space and motion prior. To this end, we leverage a cycle-consistency strategy to distill a forward deformation field from MonoNPHM, such that it becomes compatible with 3D Gaussian Splatting. We proposed an effective regularization strategy and adjust the adaptive density control strategy, both of which are important for ideal avatar quality. We show that utilizing per-Gaussian features to conditioning the complete dynamics module helps, compared to limiting the influence to appearance changes. In our experiments, we significantly outperform the previous state-of-the-art avatars on self-reenactment. Finally, we showed the applicability of our avatars beyond a controlled multi-view set-up by animating them from monocular RGB video trackings.